OpenStudy (anonymous):

If x>3, then the equation √(x+9)= x-3 equals which of the following? x=x^2 x=x^2+18 x=x^2-6x x=x^2-6x+9 x=x^2-6x+18

OpenStudy (anonymous):

If x>3, then both side are positive that means, u can square the equation

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

\[\sqrt{x+9}=x-3\] \[\sqrt{x+9}^2=(x-3)^2\]\[\ x+9=x^2-6x+9\]

OpenStudy (anonymous):

ahh, thanx!